Fe2O3 (s) + CO (g) ----> Fe (s) + CO2  

13.5 g of Fe2O3 is reacted with 15.0 g of Co. 8.30 g of Fe is obtained. Identify the limiting reagent and calculate the theoretical yield and percent yield for Fe.
Answer #1

Balanced reaction :

Fe2O3 + 3CO -> 2Fe + 3CO2

According to the reaction,

1 mole Fe2O3 require 3 mole CO

159.69 g Fe2O3 require 3×28 = 84 g CO

13.5 g Fe2O3 require 84×13.5/159.69 g CO

= 7.10 g CO

Since we have more CO available hence Fe2O3 is limiting reagent.

According to the reaction,

159.69 g Fe2O3 give 2 mole Fe

159.69 g Fe2O3 give 2×55.8 = 111.69 g Fe

13.5 g Fe2O3 give 111.69×13.5/159.69 g Fe

= 9.44 g Fe

Hence, theoretical yield = 9.44 g

Percent yield = 8.30×100/9.44

= 87.92%